Grade 12th passMechanics

2 boy's of masses 40 and 60kg are standing at the opposite ends of a boat of length 8m and mass 120 kg in river .what is the displacement of boat relative to water if boy's interchange third position

Vaishnavi
7 Years agoGrade 12th pass

1 Answer

AKSHANSH PAREEK
7 Years ago
displacement of plank with  respect to 40kg boy
use center of mass
=60*8/120+60+40
displacement of plank with  respect to 60kg boy
=40*8/120+60+40
 
net displacement of planck
=60*8/120+60+40  -  40*8/120+60+40
=7/11m
 
please please  please ….............. approve if you are satisfied